OpenStudy (anonymous):

Phoung volunteers at an SPCA that always houses 8 dogs. She notices that she uses seven 35-pound bags of dry dog food every two months. A new, larger SPCA facility that houses 20 dogs will open soon. Help Phoung estimate the amount of dry dog food that the facility should order every three months. Explain your reasoning.

OpenStudy (anonymous):

the answer is about

OpenStudy (dumbcow):

First find how much food is used every month for 8 dogs 7 bags in 2 months so 7/2 bags for 1 month ok now set up ratio to find number of bags needed every month for 20 dogs \[\frac{\frac{7}{2}}{8} = \frac{x}{20}\] crossmultiply to find x, then multiply by 3 to find food needed for 3 months for 20 dogs

OpenStudy (anonymous):

26.25

OpenStudy (dumbcow):

correct good job

OpenStudy (anonymous):

thank you so much